Suppose a $3 per-unit tax is imposed on the sellers of this good.

1) What is the effective price that sellers will receive for the good after the tax is imposed?

2) What price will buyers pay for the good after the tax is imposed?

3)How much is the burden of this tax on the buyers/sellers in this market?

How do you calculate it? Please explain.Price 20 18 16 14 12 10 8 6 4 D 10 12 14 16 uantity 4 6 rt
Answer #1

1)

Q = a - bP

b = Change in Q / Change in P

= 2/2

= 1

8 = a - b*12

8 = a - 12

a = 20

Q = 20 - P Demand equation.

Similarly supply curve:

Qs = C+bP

b = 1

8 = C + 1*8

C = 8 -8

= 0

Qs = P

Equilibrium:

20 -P = P

20 = 2P

P = 10

Q = 10

New equilibrium after Tax

20-P = P -3

23 = 2P

P = 23/2

=11.5

Seller will receive only 8.5

or it will pay 1.5 tax.

b)

Buyer will be pay 11.5 per unit.

c)

Both share equal burden of tax.

Price rise by 1.5, so buyer pays only 1.5 and rest of amount that is 1.5 would be paid by seller.